# SPIN-Bench: How Well Do LLMs Plan Strategically and Reason Socially?

### v4

@@ 3.4 Evaluation metrics
For classical planning tasks, we measure both accuracy and N-Step Look Ahead (see Appendix B.2 for the formula). In competitive games, we compare LLM moves against solver-generated top-$k$ actions and track head-to-head outcomes against the solver. We also maintain a leaderboard of internal Elo ratings (do not directly reflect performance among humans). This helps us gauge the ”skill gap” among different LLMs and compare them against human baselines. For Hanabi, we use the final score to assess cooperative efficiency. Meanwhile, Diplomacy performance is evaluated through factual consistency, order-level success rates, and final game outcomes (see Appendix B.5 for details).
@@ 4.2 Core planning competencies: factual retrieval,spatial reasoning and constraint following
In Diplomacy, we design and categorize several factual queries into one-hop (e.g., ”Which locations does Russia control?”) vs. multi-hop (e.g., ”Which territories can France attack, and how many units of support are needed at least?”) to further check models’ factual retrieval in a highly strategic environment. As summarized in Figure 2, nearly all LLMs do well on basic location or adjacency checks but degrade by a large margin on ”Attackable” and ”Attack Analysis,” which demand deeper, multi-hop inference. Again, o1 and o1-preview lead but still exhibit significant drops compared to simpler tasks. We also examined the factual retrieval performance of models within the classical planning task in Appendix C.1. Our evaluation indicates that recent reasoning models can achieve accuracy exceeding 90%, while other models exhibit significant performance degradation as problem complexity increases.
@@ 4.3.1 Complex or expansive action spaces
We further examine the results of 10 domains in classical planning tasks by analyzing two key measures: the average number of legal actions available at each step and the total state–action space, defined as the complete set of possible state–action pairs given an initial state and action space. To evaluate their impact on model performance, we calculate the correlation between accuracy and these two metrics. As illustrated in Figure 3(a) and (b), accuracy demonstrates a stronger negative correlation with the total state–action space compared to the average number of legal actions. This implies that cognitively, the model ”carries” the burden of large future branching factors, even if only a few choices are valid at each step.
@@ B.2 Classification and description of PDDL domains
    • freecell: Card puzzle variant with limited ”free cells,” each used as a temporary resource to reshuffle the card configuration.
@@ B.3 Two-player competitive games
Tic Tac Toe is a classic and simple two-player game played on a 3x3 grid. Each player takes turns marking a space on the grid, with one player using ”X” and the other ”O.” The objective is to be the first to align three of your marks horizontally, vertically, or diagonally. Despite its straightforward rules, the game challenges players to think strategically, anticipate their opponent’s moves, and block potential winning combinations. Tic Tac Toe is widely popular because it is easy to learn, quick to play, and requires minimal setup. So in our paper, we use Tic Tac Toe as the simplest setting testing LLMs’ ability to understand basic rules and strategy.
@@
For the step-wise evaluation metrics, we use the scores given by the solver and Stockfish, among all the trajectories in LLM-vs-Solver setting. For Tic Tac Toe and Connect Four, use the solver to generate scores for each move of LLM at each state. For Chess, we use Stockfish level 20 as the rating machine for step-wise analysis, on each step in the game between LLMs and Stockfish level 0. We use ”centipawns” as a measurement of the advantage of a chess board. A centipawn is equal to 1/100 of a pawn. These values are essentials in computer chess to evaluate positions. For each pair of the game, we conduct 4 repetitive experiments.
@@ B.4 Multi-player cooperative game - Hanabi
    Give Information: The player points out the cards of either a given number or a given color in the hand of another player (examples: ”This card is your only red card,” ”These two cards are your only 3s”). The information given must be complete and correct. Giving information consumes one information token.
@@ B.5.1 Game introduction
Diplomacy proceeds by seasons, beginning in the year 1901, with each year divided into two main seasons: the ”Spring” and ”Fall” (Autumn) moves. Each season is further divided into negotiation and movement phases, followed by ”retreat” or ”disband” adjustments and an end-of-the-year Winter phase of new builds or removals following the Fall adjustments.
@@
After the negotiation period, players write secret orders for each unit; these orders are revealed and executed simultaneously. A unit can move from its location to an adjacent space, support an adjacent unit to hold an area in the event of an attack, support another unit to attack a space into which it could move itself, or hold defensively. In addition, fleets may transport armies from one coast space to another when in a chain called a ”convoy”. Only one unit may occupy each region. If multiple units are ordered to move to the same region, only the unit with the most support moves there. If two or more units have the same highest support, a standoff occurs and no units ordered to that region move. A unit ordered to give support that is attacked has those orders canceled and is forced to hold, except in the case that support is being given to a unit invading the region from which the attack originated (in which case the unit that had been ordered to give support must retreat from, rather than hold, its position). Certain spaces on the board have two coasts and here a player must specify which one they want their fleet to occupy. A fleet can only move to coasts and oceans that border the coast that it is on. For example, a fleet occupying the southern coast of Bulgaria cannot move into Romania or the Black Sea, but a fleet on the east coast could.
@@ B.5.2 Evaluation metrics
Diplomacy is a highly intricate strategy game, so to effectively assess the basic proficiency of LLMs in playing it, we designed this basic skill(BS) evaluation. In this experiment, a single LLM is assigned the role of France, while the remaining six major powers - England, Germany, Italy, Austria, Russia, and Turkey are controlled by neutral agents. These neutral powers adopt a passive stance, not doing movement or building, while only disbanding their units when necessary. We also disable the negotiation part in each movement phase, to check whether LLM has the ability to play Diplomacy. We use this “all‐neutral” configuration purely in our first‐round evaluation to isolate each model’s basic game proficiency. By setting all other powers ”neutral” will make the game super simple and easy. If a player can understand the game rules and basic actions, they can win the game easily. Once we identify models that demonstrate reliable core mechanics under these simplified conditions, we advance them to a second evaluation stage involving full multiplayer scenarios where social reasoning, negotiation dynamics, and long‐term planning become critical.
@@ C.1 PDDL
Here, we investigate whether LLMs can reliably retrieve key facts from a planning trajectory. For each task, we provide the model with the initial state ($s_{\text{ini}}$), the goal state ($S_{G}$), and trajectories that explicitly detail the actions along with their corresponding state transitions. We control the number of mutable properties; for example, in a task involving moving balls, the ball location is considered mutable, whereas ball color might also be a property—but if no action can alter this property, we do not count it as mutable. We then prompt the model to report the state at a specified step. Figure 6 illustrates how retrieval accuracy varies with trajectory length. Notably, o1 performs most consistently, confirming that it ”reads” multi-step expansions more accurately than either GPT-4o or o1-mini.
@@ C.3.1 Basic skill evaluation
Table 9: Performance of LLMs and human in the single-player setting of Diplomacy. ”-” means that game was forcibly terminated after 20 years in the game, indicating that model can not win the game in that time.
@@ Appendix D Discussion/limitation
D.1 Limited complixty on ”social Intelligence”
@@ E.2 Cooperative game: Hanabi
{"reason": string, "action": int} And the action should be one of the legal actions provided above.
@@
"reason": "Card 5 in your hand is known to be B2, which is exactly the next playable card for the blue stack (currently at B1). Playing it will advance the blue firework without risk.",
"action": "9"
@@ 5.2. You should take control more and more supply centers
    Possible Orders: Diplomacy features an extensive action space with numerous legal orders available in each phase. To reduce the complexity for the agent in adhering to constraints, we supply all legal orders for each of the agent’s units. These orders are presented as a list, with each element being a dictionary in the following format: {”TURKEY”: [{’SMY’: [’F SMY - SYR’, ’F SMY S A CON’, …]}, …]}. In this structure, the key denotes the power being controlled by that agent, and the value is a list of orderable units along with their corresponding possible orders.

### v5

@@ 3.4 Evaluation metrics
For classical planning tasks, we measure both accuracy and N-Step Look Ahead (see Appendix B.2 for the formula). In competitive games, we compare LLM moves against solver-generated top-$k$ actions and track head-to-head outcomes against the solver. We also maintain a leaderboard of internal Elo ratings (do not directly reflect performance among humans). This helps us gauge the "skill gap" among different LLMs and compare them against human baselines. For Hanabi, we use the final score to assess cooperative efficiency. Meanwhile, Diplomacy performance is evaluated through factual consistency, order-level success rates, and final game outcomes (see Appendix B.5 for details).
@@ 4.2 Core planning competencies: factual retrieval,spatial reasoning and constraint following
In Diplomacy, we design and categorize several factual queries into one-hop (e.g., "Which locations does Russia control?") vs. multi-hop (e.g., "Which territories can France attack, and how many units of support are needed at least?") to further check models’ factual retrieval in a highly strategic environment. As summarized in Figure 2, nearly all LLMs do well on basic location or adjacency checks but degrade by a large margin on "Attackable" and "Attack Analysis," which demand deeper, multi-hop inference. Again, o1 and o1-preview lead but still exhibit significant drops compared to simpler tasks. We also examined the factual retrieval performance of models within the classical planning task in Appendix C.1. Our evaluation indicates that recent reasoning models can achieve accuracy exceeding 90%, while other models exhibit significant performance degradation as problem complexity increases.
@@ 4.3.1 Complex or expansive action spaces
We further examine the results of 10 domains in classical planning tasks by analyzing two key measures: the average number of legal actions available at each step and the total state–action space, defined as the complete set of possible state–action pairs given an initial state and action space. To evaluate their impact on model performance, we calculate the correlation between accuracy and these two metrics. As illustrated in Figure 3(a) and (b), accuracy demonstrates a stronger negative correlation with the total state–action space compared to the average number of legal actions. This implies that cognitively, the model "carries" the burden of large future branching factors, even if only a few choices are valid at each step.
@@ 6 Conclusion
Acknowledgments
We gratefully acknowledge the Sentient Foundation for sponsoring compute credits, and we thank the technical staff at Sentient for their valuable discussions during this work.
@@ B.2 Classification and description of PDDL domains
    • freecell: Card puzzle variant with limited "free cells," each used as a temporary resource to reshuffle the card configuration.
@@ B.3 Two-player competitive games
Tic Tac Toe is a classic and simple two-player game played on a 3x3 grid. Each player takes turns marking a space on the grid, with one player using "X" and the other "O." The objective is to be the first to align three of your marks horizontally, vertically, or diagonally. Despite its straightforward rules, the game challenges players to think strategically, anticipate their opponent’s moves, and block potential winning combinations. Tic Tac Toe is widely popular because it is easy to learn, quick to play, and requires minimal setup. So in our paper, we use Tic Tac Toe as the simplest setting testing LLMs’ ability to understand basic rules and strategy.
@@
For the step-wise evaluation metrics, we use the scores given by the solver and Stockfish, among all the trajectories in LLM-vs-Solver setting. For Tic Tac Toe and Connect Four, use the solver to generate scores for each move of LLM at each state. For Chess, we use Stockfish level 20 as the rating machine for step-wise analysis, on each step in the game between LLMs and Stockfish level 0. We use "centipawns" as a measurement of the advantage of a chess board. A centipawn is equal to 1/100 of a pawn. These values are essentials in computer chess to evaluate positions. For each pair of the game, we conduct 4 repetitive experiments.
@@ B.4 Multi-player cooperative game - Hanabi
    Give Information: The player points out the cards of either a given number or a given color in the hand of another player (examples: "This card is your only red card," "These two cards are your only 3s"). The information given must be complete and correct. Giving information consumes one information token.
@@ B.5.1 Game introduction
Diplomacy proceeds by seasons, beginning in the year 1901, with each year divided into two main seasons: the "Spring" and "Fall" (Autumn) moves. Each season is further divided into negotiation and movement phases, followed by "retreat" or "disband" adjustments and an end-of-the-year Winter phase of new builds or removals following the Fall adjustments.
@@
After the negotiation period, players write secret orders for each unit; these orders are revealed and executed simultaneously. A unit can move from its location to an adjacent space, support an adjacent unit to hold an area in the event of an attack, support another unit to attack a space into which it could move itself, or hold defensively. In addition, fleets may transport armies from one coast space to another when in a chain called a "convoy". Only one unit may occupy each region. If multiple units are ordered to move to the same region, only the unit with the most support moves there. If two or more units have the same highest support, a standoff occurs and no units ordered to that region move. A unit ordered to give support that is attacked has those orders canceled and is forced to hold, except in the case that support is being given to a unit invading the region from which the attack originated (in which case the unit that had been ordered to give support must retreat from, rather than hold, its position). Certain spaces on the board have two coasts and here a player must specify which one they want their fleet to occupy. A fleet can only move to coasts and oceans that border the coast that it is on. For example, a fleet occupying the southern coast of Bulgaria cannot move into Romania or the Black Sea, but a fleet on the east coast could.
@@ B.5.2 Evaluation metrics
Diplomacy is a highly intricate strategy game, so to effectively assess the basic proficiency of LLMs in playing it, we designed this basic skill(BS) evaluation. In this experiment, a single LLM is assigned the role of France, while the remaining six major powers - England, Germany, Italy, Austria, Russia, and Turkey are controlled by neutral agents. These neutral powers adopt a passive stance, not doing movement or building, while only disbanding their units when necessary. We also disable the negotiation part in each movement phase, to check whether LLM has the ability to play Diplomacy. We use this “all‐neutral” configuration purely in our first‐round evaluation to isolate each model’s basic game proficiency. By setting all other powers "neutral" will make the game super simple and easy. If a player can understand the game rules and basic actions, they can win the game easily. Once we identify models that demonstrate reliable core mechanics under these simplified conditions, we advance them to a second evaluation stage involving full multiplayer scenarios where social reasoning, negotiation dynamics, and long‐term planning become critical.
@@ C.1 PDDL
Here, we investigate whether LLMs can reliably retrieve key facts from a planning trajectory. For each task, we provide the model with the initial state ($s_{\text{ini}}$), the goal state ($S_{G}$), and trajectories that explicitly detail the actions along with their corresponding state transitions. We control the number of mutable properties; for example, in a task involving moving balls, the ball location is considered mutable, whereas ball color might also be a property—but if no action can alter this property, we do not count it as mutable. We then prompt the model to report the state at a specified step. Figure 6 illustrates how retrieval accuracy varies with trajectory length. Notably, o1 performs most consistently, confirming that it "reads" multi-step expansions more accurately than either GPT-4o or o1-mini.
@@ C.3.1 Basic skill evaluation
Table 9: Performance of LLMs and human in the single-player setting of Diplomacy. "-" means that game was forcibly terminated after 20 years in the game, indicating that model can not win the game in that time.
@@ Appendix D Discussion/limitation
D.1 Limited complixty on "social Intelligence"
@@ E.2 Cooperative game: Hanabi
{reason: string, action: int} And the action should be one of the legal actions provided above.
@@
reason: Card 5 in your hand is known to be B2, which is exactly the next playable card for the blue stack (currently at B1). Playing it will advance the blue firework without risk.,
action: 9
@@ 5.2. You should take control more and more supply centers
    Possible Orders: Diplomacy features an extensive action space with numerous legal orders available in each phase. To reduce the complexity for the agent in adhering to constraints, we supply all legal orders for each of the agent’s units. These orders are presented as a list, with each element being a dictionary in the following format: {"TURKEY": [{’SMY’: [’F SMY - SYR’, ’F SMY S A CON’, …]}, …]}. In this structure, the key denotes the power being controlled by that agent, and the value is a list of orderable units along with their corresponding possible orders.